OpenStudy (hkmiu):

Explain how to write the series 2 + 12 +72 + 432 + 2592 using sigma notation.

OpenStudy (akashdeepdeb):

We need to find the sum of n terms right! So find the nth term first = ar^(n-1) = 2*6^(5-1) 2*6^4 So n varies from 1 to 5 So in sigma notation it'll be \[\sum_{1}^{5} 2*6^(n-1)\]

OpenStudy (akashdeepdeb):

Becuase it is a geometric progression! :)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (hkmiu):

Thank you @AkashdeepDeb !

OpenStudy (hkmiu):

I've been stuck on this for awhile, haha.

OpenStudy (akashdeepdeb):

:)